OpenStudy (anonymous):

find the solutions to the systems on nonlinear equations y=3x x^2+y^2=40

OpenStudy (anonymous):

\[\begin{cases}y=3x\\x^2+y^2=40\end{cases}\] Plug the first equation into the second: \[x^2+(3x)^2=40\\ x^2+9x^2=40\\ 10x^2=40\\ x^2=4\\ x=\pm2\] Since \(y=3x\), that means \(y=3\times2=6\) and \(y=3\times(-2)=-6\). Your solutions are then \((2,6)\) and \((-2,-6)\).